OpenStudy (senia01):

A tree that is 8 feet tall is growing at a rate of one foot each year. A different tree that is 10 feet tall is growing at a rate of 1/2 foot each year. 1.Write expressions to show how tall each tree will be in x years. 2. Use your expressions to find how many years it will take for the trees to be the same height.

OpenStudy (neonumbrella5115):

Have you tried to write an expression? Try using y=mx+b.

OpenStudy (senia01):

i did y=8+1x and y=10+.5x but i didnt think it was right

OpenStudy (senia01):

@neonumbrella5115

OpenStudy (neonumbrella5115):

Those are correct! So you've done step 1, now you need to find when the equations will be equal to each other. Using substitution, substitute the y from one equation for the y of another to get: 1x+8=0.5x+10 Then solve. Care you good with solving?

OpenStudy (senia01):

I got 4

OpenStudy (neonumbrella5115):

That's correct good job! Make sure to include the label of 4 years. If you need any more help just tag me!